OpenStudy (anonymous):

If a stone is thrown vertically upward from the surface of the moon with a velocity of 10 m/s, its height (in meters) after t seconds is h(t)=10t−0.83t^2. (a) What is the velocity of the stone after 8 seconds? (b) What is the velocity of the stone after it has risen 25 m?

OpenStudy (anonymous):

I already figured out part a. I just need assistance with b! Please help!

OpenStudy (anonymous):

Omg. It's supposed to say h(t)=10t-0.83t^2.

OpenStudy (badhi):

since the equations on constant acceleration stands, your given expression is similar to $$S=ut +\frac{1}{2}at^2$$ here $$u=10,\quad a=0.83\times 2=1.66$$ to find the velocity at 25m use $$v^2=u^2+2as$$